OpenStudy (erinweeks):

A mathematics journal has accepted 14 articles for publication. However, due to budgetary restraints only 9 articles can be published this month. How many ways can the journal editor assemble 9 of the 14 articles for publication?

OpenStudy (erinweeks):

A. 726,485,760 B. 14 C. 2,002 D. 126

OpenStudy (anonymous):

you are being asked for how many ways to choose 9 out of a set of 14 in math this is called "14 choose 9" the notation is \(\dbinom{14}{9}\) and it is calculated via \[\dbinom{14}{9}=\frac{14\times 13\times 12\times 11\times 10}{5\times 4\times 3\times 2}\] cancel first, multiply last you get \[14\times 13\times 11=2002\]

OpenStudy (anonymous):

many calculators have a key for that, or you can use this http://www.wolframalpha.com/input/?i=14+choose+9

OpenStudy (erinweeks):

well 2,002 is wrong :/i already used that answer and i got it wrong

OpenStudy (mathmate):

The question is not clear. Assemble probably is meant to say order is important, which means the answer should be 14P9, or 726485760.

jimthompson5910 (jim_thompson5910):

If order matters, then n P r = (n!)/(n-r)! 14 P 9 = (14!)/(14-9)! 14 P 9 = (14!)/(5)! 14 P 9 = (14*13*12*11*10*9*8*7*6*5!)/(5)! 14 P 9 = 14*13*12*11*10*9*8*7*6 14 P 9 = 726,485,760 So if order matters, then there are 726,485,760 different ways

OpenStudy (mathmate):

14P9 = 14!/(14-9!